# PROFIT: A Specialized Optimizer for Deep Fine Tuning

### v2

@@ 4.2 Image Classification
Full [21] 68.9 87.7 64.3 87.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1
Linear [21] 64.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.5 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
@@
Table 5: Image classification accuracy on each task in the VTAB-1k benchmark. We compare PROFIT against standard fine-tuning strategies, including full fine-tuning and linear fine-tuning (denoted by Full [18] and Linear [18]). Results are reported across various datasets categorized into Natural, Specialized, and Structured domains. A naive implementation of PROFIT performs suboptimally due to the violation of the proximity condition in this setting, but when supplemented with AdamW warm-up PROFIT (warm-up) demonstrates superior performance and consistently achieves higher accuracy scores in multiple datasets. For each block (Natural, Specialized, and Structured), the highest accuracy are highlighted in bold.
Next, we demonstrate the effectiveness of PROFIT on the image classification task. CIFAR10 and CIFAR100 [28] are the de-facto benchmarks for image classification. The dataset consists of $32\times 32$ input images, and the task for the network is to classify an image into one of $K$ categories.
We first train a network on CIFAR10 (pre-training), and fine-tune the network on CIFAR100. As discussed earlier, one of the assumptions for PROFIT is that of proximality, where both the pretraining and fine-tuning datasets need to be from a similar domain distribution. This assumption is met for our choice of CIFAR10 and CIFAR100, as they are both labeled subsets of Tiny Images [49].
We experiment with various choices of backbones, including ResNet-18 [17], ViT-Tiny, and ViT-Small [12]. We compare PROFIT against a variety of the baselines, such as Stochastic Gradient Descent (SGD), Adam [23], Lookahead [58], and Adan [54]. We run the experiments 3 times and report the best results, but see low variance in our results.
Our results are shown in Table 2, which show that PROFIT is able to perform better on the fine-tuning task across all backbone networks and optimizers. We also show better CIFAR10 accuracy after fine-tuning, indicating that our method is better at maintaining old task performance as well. For example, fine-tuning a ResNet-18 with Adan [54] optimizer leads to 35.17% accuracy on CIFAR10 and 72.70% accuracy on CIFAR100, while using PROFIT shows better original task performance (35.26% accuracy) and fine-tuning task performance (74.70% accuracy). With ViT-Tiny, Adam fine-tuning results in 55.64% accuracy on CIFAR10 and 61.35% accuracy on CIFAR100, but our method comfortable outperforms with 58.53% on CIFAR10 and 62.20% on CIFAR100. fine-tuning ViT-Small using Adam shows 58.60% accuracy on CIFAR10 and 63.93% accuracy on CIFAR100, however, PROFIT shows better performance with 59.02% accuracy on CIFAR10 and 65.44% accuracy on CIFAR100.
We ablate $n_{\text{ref}}$ in Table 4. In general, we observe that increasing $n_{\text{ref}}$ leads to better performance on the original task while compromising performance on the fine-tuning task. This follows from our discussion in Sec 3.2, since $n_{\text{ref}}$ controls the degree of exploration away from the stable point. Nonetheless, we recommend $n_{\text{ref}}=1$ as a starting point.
We also ablate performance on $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ in Table 4, which shows the best choice may vary for the choice of backbone. In general, larger values of $\lambda_{\text{ref}}$ promotes new task accuracy, while smaller values mitigate catastrophic forgetting effectively. These results are reasonable, as smaller values of $\lambda_{\text{ref}}$ equates to the reference point lying closer to the original model state. However, we always were able to beat the baseline on both old and new task accuracies simultaneously.
We also use this setting to substantiate Assumption 2 (Sec 3.4), which states PROFIT does not work well without a converged model. When we train a ResNet-18 from scratch on CIFAR100 using PROFIT, we get 1.05% accuracy, which is as good as random guessing.
4.3 Visual Task Adaptation Benchmark
The Visual Task Adapation Benchmark (VTAB-1K) [56] is a popular representation learning benchmark to evaluate generalization across a diverse set of classification tasks. It consists of 19 different tasks from a broad spectrum of domains, with the aim to learn feature representations which perform well on all of these tasks. Performance is measured by computing classification accuracy on each of these 19 tasks.
We use ViT-Base [20] with ImageNet pretraining as the backbone for our experiments. This setting is an example of a setting where the original (ImageNet) and fine-tuning distributions (VTAB-1K) significantly differ, violating Assumption 1 (Sec 3.4). While our method is not well tuned for these fine-tuning settings, this situation is common in practice. In this section, we show how a practitioner can leverage our method in such settings.
We show our results in Table 5, and compare our method with AdamW [34] fine-tuning of the entire model (Full – row 1) as well as final layer fine-tuning (Linear – row 2). Full fine-tuning using PROFIT does not work well, showing poor performance across all 19 tasks in the benchmark.
However, we provide a training recipe on how PROFIT can be used in such settings. First, we fine-tune the model towards the target distribution with an AdamW warmup for 10 epochs, and then apply PROFIT to fine-tune for the remaining 90 more epochs (PROFIT(warmup)). Intuitively, this moves the model’s distribution somewhere between the pre-training and fine-tuning distribution, which makes it amenable to the structure of PROFIT. This approach outperforms full model fine-tuning with AdamW. So while our method as expected is not recommended for non-proximal settings (as discussed in Section 3.4), using a short warmup phase with another optimizer allows PROFIT to substantially outperform full fine-tuning with standard optimizers.
4.4 Large-Scale Robotics Motion Prediction
We now show how PROFIT performs on an extremely challenging benchmark with a large, high-dimensional dataset. The Waymo Open Motion Dataset (WOMD)  [13] is a large-scale driving dataset collected from realistic scenarios. The task is to predict future trajectories of an agent over the next 8 seconds, given multi-modal observations in the last second, including the agent’s history, nearby agent histories, map information, and traffic light states. The prediction model follows a state-of-the-art early-fusion transformer, Wayformer [38], by fusing multi-modal input features through a self-attention transformer and predicting future trajectory samples using learned latent queries.
We first train a model on car trajectory prediction, and then fine-tune the model on data from the same class (car) as well as different classes (pedestrian). For our baselines, we use the AdamW [34] optimizer. WOMD offers an important difference from previous settings in that we can see how well PROFIT performs on (1) fine-tuning a model on the same data, and (2) fine-tuning a model on a different domain-shifted task. For instance, CIFAR100 is explicitly part of the same distribution as CIFAR10, but pedestrian trajectories are semantically different than car trajectories.
Figure 3: Results for PROFIT on Waymo Open Motion Dataset. (a) Error curves through training tabulated for FDE at 3s and 8s (top) and ADE at 3s and 8s (bottom). PROFIT outperforms both fine-tuning baselines by a sizable margin. (b) Visualizations of motion prediction outputs for both the baseline fine-tune model (top) and PROFIT (bottom). Trajectory ground truth is shown as a shaded bar and denser lines represent more confident predictions. Although the differences are often subtle, PROFIT generally produces more confidently correct predictions. Best viewed in color.
Method Target Class ADE@3s (m) ADE@8s (m) FDE@3s FDE@8s
Baseline - 0.461 1.327 1.024 2.581
fine-tune (F) Car 0.458 1.322 1.021 2.548
fine-tune (H) Car 0.456 1.303 1.009 2.507
PROFIT Car 0.454 1.299 1.008 2.489
fine-tune (F) Ped 0.214 0.621 0.465 1.242
fine-tune (H) Ped 0.232 0.724 0.508 1.544
PROFIT Ped 0.203 0.579 0.427 1.145
Figure 4: We compare the baseline with PROFIT on an example from DriveLM. The model fine-tuned with PROFIT is able to perceive the traffic light and black sedan in the scene, while the baseline (fine-tuned with AdamW) does not detect the traffic light and hallucinates the presence of a white truck. Consequently, the baseline suggests running the red light, while our method follows traffic rules by staying stationary. Best view in color.
The results are shown in Figure 3 and Table 6. Average Distance Error (ADE) measures the average distance between ground truth and prediction at each point of the predicted trajectories, while Final Distance Error (FDE) only considers the error in the final point of the predicted trajectories. We see that in all cases, PROFIT outperforms the baseline, with sizable benefits in the car-to-pedestrian fine-tuning task. Consistent with the results in other settings, fine-tuning with head only provides subpar performance on the car-to-pedestrian task, while full fine-tuning does reasonably well but is still outstripped by PROFIT. We conclude that PROFIT leads to a better results to re-purpose a model for a new, correlated setting.
Backbone Optimizer Accuracy (%) $\uparrow$ GPT Score (%) $\uparrow$ Match Score (%) $\uparrow$ BLEU-4 $\uparrow$ ROUGE_L $\uparrow$ CIDEr $\uparrow$ Final Score $\uparrow$
LLAMA-Adapter-v2 [15] AdamW 62.21 70.57 36.33 0.541 7.16 0.023 56.98
PROFIT 67.88 72.12 37.37 0.557 7.28 0.035 59.16
Table 7: Results on the DriveLM Benchmark. We fine-tune LLAMA-Adapter-v2 [15] and compare PROFIT with AdamW, the de-facto method for fine-tuning VLMs. Our method shows noticeable improvements on all metrics, showcasing the applicability of PROFIT to large foundational models.
Backbone $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$ Accuracy (%) $\uparrow$ GPT Score (%) $\uparrow$ Match Score (%) $\uparrow$ BLEU-4 $\uparrow$ ROUGE_L $\uparrow$ CIDEr $\uparrow$ Final Score $\uparrow$
LLAMA-Adapter-v2 [15] 10 65.82 71.99 37.21 0.550 7.21 0.029 58.54
100 65.83 72.19 37.16 0.550 7.22 0.029 58.64
1000 65.74 72.07 37.02 0.550 7.22 0.030 58.54
10000 67.88 72.12 37.37 0.557 7.28 0.035 59.16
We also see minor but noticeable improvements in the car-to-car benchmark, suggesting that PROFIT can be used to extract more performance from any model that has already converged. We could even imagine a scenario in which regular additional training runs using PROFIT are enabled as a standard model maintenance practice.
4.5 Driving with Vision Language Models
DriveLM [46] is a visual question-answering (VQA) benchmark to evaluate vision-language models (VLMs) for an autonomous driving setting. The DriveLM-nuscenes dataset contains sensor inputs and a text prompt, and the network aims to provide accurate textual responses for prompts which probe perception, prediction, and planning capabilities respectively.
The DriveLM benchmark leverages a variety of VQA metrics for evaluation. BLEU [39] evaluates precision by measuring n-gram similarity between generate and reference texts. ROUGE_L [31] evaluates recall by measuring the longest common subsequence between the generated and reference texts. CIDEr [51] evaluates quality by computing the cosine similarity between the n-gram TF-IDF vectors for the generated and reference sentences. GPT Score aims to captures semantic nuances missed by aforementioned metrics by using ChatGPT (GPT-3.5-Turbo) as an evaluator. In addition to these metrics, perception accuracy is evaluated using the ground-truth objects in the scene, while prediction accuracy is evaluated over discretized future states. Match score evaluates whether the VLM correctly understands the order in which to attend to other agents in the scene. The final score for this benchmark is a weighted average of these metrics. For further details, we refer the reader to the evaluation criteria for the DriveLM Challenge [11].
We use a pre-trained LLAMA-Adapter-v2 [15], which performs bias tuning on LLAMA-7B [50]. We compare PROFIT with AdamW [34], the de-facto method for fine-tuning VLMs.
We report our results on the validation set of DriveLM in Table 7. PROFIT improves perception and prediction accuracy over AdamW by 5.6%, GPT score by 1.5%, Match Score by 1%, and the Final Score by 2. In addition, PROFIT provides better outputs for VQA, as seen by improved performance on language metrics (BLEU-4, ROUGE_L, CIDEr). We provide an inference visualization of our method in Fig 4. The baseline, fine-tuned with AdamW, is unable to detect the traffic light and black sedan in the scene. This leads to the suggested plan of running the red light, which is a potentially catastrophic error. On the other hand, PROFIT successfully identifies the red light and follows traffic regulations by planning to remain stationary. We conclude PROFIT is a useful tool in the era of VLMs to extract better performance.
We ablate performance on $\frac{\lambda_{\text{main}}}{\lambda_{\text{ref}}}$, a key hyper-parameter to PROFIT, in Table 8. In general, larger values encourage exploration and improves fine-tuning task accuracy. This follows Sec 4.2, since smaller ${\lambda_{\text{ref}}}$ implies the reference step away from the pre-trained model is smaller.
5 Conclusion
We propose PROFIT, an optimizer which is designed to robustify models during fine-tuning. Unlike standard optimizers which cannot assume much about a training setting, our ability to assume confidence about the prior state of a model allows PROFIT to act as an effective regularizer and prevent a model’s weights from diverging too much from its previously good state. We showed that PROFIT performs well in various settings: (1) fine-tuning to new data off the data manifold in our toy example, (2) fine-tuning to new CIFAR100 data from CIFAR10 pretraining, (3) fine-tuning to a new data distribution in VTAB-1K, and (4) fine-tuning a model on a completely new task as well as on the same exact task on large-scale motion prediction. In all cases, PROFIT surpassed the performance of the near-ubiquitously used, standard fine-tuning approaches, while being just as easy to integrate within a generic model training pipeline. We believe that PROFIT should not only be an important tool in the deep learning practitioner’s arsenal but opens the discussion around developing a new class of optimizers that will crucially support the new age of deep learning where fine-tuning becomes the primary training paradigm.
6 Acknowledgment
We thank Carl Vondrick, David Hayden, Mao Ye, Navaneeth Bodla, Amirreza Shaban, Greg Meyer, and Sidd Srinivasa for their guidance and discussions which help shape our work.

### v3

@@ C.5.1 Implementation Details
Full Jia et al. [2022] 68.9 87.7 64.3 87.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1
Linear Jia et al. [2022] 64.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.5 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2
VPT Jia et al. [2022] 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8
LoRA Hu et al. [2022] 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.0
NOAH Zhang et al. [2025] 69.6 92.7 70.2 99.1 90.4 86.1 53.7 84.4 95.4 83.9 75.8 82.8 68.9 49.9 81.7 81.8 48.3 32.8 44.2
@@
Table 13: Image classification accuracy on VTAB-1K. Bold = best, Underline = second-best. Ties for best are all bolded; no second-best is shown in such cases. We compare PROFIT against standard fine-tuning strategies, including full fine-tuning and linear fine-tuning (denoted by Full and Linear). A naive PROFIT violates the proximity condition and performs poorly, but with AdamW warm-up it achieves consistently higher accuracy.
Appendix D Limitations
As discussed in Sec 3.2, PROFIT requires inference on $n_{\text{ref}}+1$ batches (2 for $n_{\text{ref}}=1$) per iteration, hence being slower than standard optimizers for fine-tuning. In practice, we observe that PROFIT converges earlier than corresponding baselines (which are trained longer for fair comparison), which may mitigate this. We also note that typical fine-tuning settings are shorter and not as involved as training from scratch, which may further alleviate this concern. Another limitation is that our method would consume slightly more memory as a result of instantiating two optimizers ($\mathbf{O}^{\text{(ref)}}$ and $\mathbf{O}$).
D.1 Memory Profiling
In training generative models, memory consumption is a bottleneck, and as mentioned in 3.4 the vanilla version of PROFIT induces some training-time memory overhead. For completeness we include those memory profiling experiments here. As model parameter counts and activation footprints grow, so does the GPU / VRAM requirement, often leading to training failures, batch-size reductions, and suboptimal hardware utilization. Memory profiling plays a critical role in assessing whether a proposed method remains practical under realistic resource constraints. In particular, when proposing a new optimizer like PROFIT, we must ensure that any additional memory overhead is manageable and that iteration time remains comparable. To validate the generality of our method, we profiled a diverse set of architectures that included both language and generative models.
We conducted a series of memory profiling experiments to compare PROFIT with the standard AdamW optimizer in a range of widely used architectures. These models were selected to represent different domains of interest for future research in fine-tuning large models. Table 14 summarizes the peak memory consumption and the per-iteration time for each setup.
Model Peak Mem. (MB) — AdamW Peak Mem. (MB) — PROFIT Iter. Time (ms) — AdamW Iter. Time (ms) — PROFIT
OPT-1.3B Zhang et al. [2022] 5532.5 6324.0 106.6 115.7
OPT-2.7B Zhang et al. [2022] 8602.4 10076.0 164.2 174.3
LLaMA3-8B Dubey et al. [2024] 20341.3 25693.4 474.7 490.4
GPT-OSS-20B Agarwal et al. [2025] 14502.5 17992.0 117.7 139.9
Stable Diffusion v1-5 Rombach et al. [2022] 5603.5 7134.5 130.3 142.9
Table 14: Peak memory usage and iteration time for AdamW vs. PROFIT. PROFIT shows modest additional memory overhead with comparable iteration latency. Note that we did not fine-tune the models with PROFIT for the accuracy here, and the experiments are for memory profiling only.
For OPT (Zhang et al. [2022]), LLaMA (Dubey et al. [2024]), and Stable Diffusion v1-5 (Rombach et al. [2022]), we adopt QLoRA (Dettmers et al. [2023]) for fine-tuning, while LoRA (Hu et al. [2022]) is used for GPT-OSS-20B. This profiling highlights PROFIT’s generality and efficiency across both language and vision–language models. As shown in Table 14, we have conducted the memory profiling experiments. The experiment shows that PROFIT is on par with AdamW, with an additional memory increase. This is because we treat both main and reference optimizers completely separately in order to better explore the hyperparameter space. As we discussed in the Section 3.4, it is expected that the memory overhead of the model training is increased approximately 25% over the AdamW (Loshchilov and Hutter [2019]) baseline.
Impact Statement
The methods presented in this work represent general advancements in the field of machine learning, and are intended to be applicable in a wide range of machine learning applications. We hope that this work will help in preventing catastrophic forgetting and overfitting to potentially harmful data, which may be a positive ethical/societal effect. Other than that consideration, although there may be broader impact considerations in any downstream applications using our work, we feel that this work taken in isolation does not engender any additional societal impacts worth noting here.